I am having trouble doing these problems. Can you help me.
2. Solve the following:
i. -2x = 3x + 4
ii. 3x/4 = 6
iii. y/6 + 1 = 9
iv. 6 = -2x/4
v. Find f(1) for f(x) = 4x3 - 3x2 - x + 2
vi. A function gives the value of C as 2 × (22/7) × r. Find C when r = 21 cm and r = 84 cm.

Answer by checkley77(12844)

i. -2x = 3x + 4
-2x-3x=4
-5x=4
x=4/-5
x=-4/5 Answer.
ii. 3x/4 = 6
3x=4*6
3x=24
x=24/3
x=8 Answer.
iii. y/6 + 1 = 9
y/6=9-1
y/6=8
y=8*6
y=48 Answer.
iv. 6 = -2x/4
6=-2x/4
-2x=4*6
-2x=24
x=24/-2
x=-12
v. Find f(1) for f(x) = 4x3 - 3x2 - x + 2
f(1)=4*1^3-3*1^2+2
f(1)=4*1-3+2
f(1)=4-3+2
f(1)=3 Answer.
vi. A function gives the value of C as 2 × (22/7) × r. Find C when r = 21 cm and r = 84 cm.
c=2(22/7)r
c=2*22/7*21
c=2*22*21/7
c=924/7
c=132 Answer.
c=2(22/7)84
c=2(22*84/7)
c=2(1848/7)
c=3696/7
c=528 Answer.